# Structure of heavy quarkonia under strong magnetic field

### v1

@@ I Introduction
Recent theoretical advances have enabled comprehensive studies of these phenomena through multiple approaches. Lattice QCD now provides first-principles calculations of hadron properties in electromagnetic fields [6], while model-based frameworks offer complementary perspectives. This vibrant research area is evidenced by numerous recent reviews covering various aspects of quarkonia in extreme environments [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17].
@@
The study of heavy hadrons containing charm or bottom quarks under magnetic fields has been approached through several theoretical frameworks. Constituent quark models [18, 19, 20, 21, 22, 23, 24] and QCD sum rules [25, 26, 27, 28] have provided valuable insights, complemented by holographic QCD approaches [29, 30, 31, 32, 33] and the complex-potential picture in thermal medium [34, 35, 36, 37, 38]. Lattice QCD simulations have further advanced our understanding by investigating the heavy quark-antiquark potential in magnetic backgrounds [39, 40, 41]. Beyond heavy quark systems, significant attention has been devoted to understanding magnetic field effects on lighter hadrons. For light mesons, studies (e.g., [42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57]) have revealed characteristic modifications, while light baryons have been examined through several approaches [58, 59, 60, 61, 62, 63, 64, 65, 66].
@@
The constituent quark model remains a powerful framework for understanding static hadron properties, as detailed in a recent review [15]. Magnetic fields influence hadrons through three main mechanisms: (i) spin mixing via the $-\bm{\mu}_{i}\cdot\bm{B}$ interaction, (ii) quark kinetic energy modifications induced by the $\bm{B}\times\bm{r}$ term, and (iii) anisotropic deformation of inter-quark potentials [67, 68, 69, 70], where lattice QCD provides crucial constraints [39, 40, 41]. Previous investigations of quarkonia [21, 22] have systematically examined these effects, particularly spin mixing and energy shifts, employing potential models. These studies utilize the cylindrical Gaussian expansion method (CGEM) to solve the two-body Schrödinger equation, an adaptation of the conventional Gaussian expansion method (GEM) [71, 72] specifically designed for systems with cylindrical symmetry under magnetic fields.
@@
Building upon previous studies [22, 21], we aim to study light-front wave functions (LFWFs) of heavy quarkonia from nonrelativistic quark model solutions under strong magnetic fields and analyze their structural modifications. Unlike the direct light-front Hamiltonian diagonalization method proposed in Ref. [24], our approach establishes a connection between magnetically-modified LFWFs and the rest-frame wave functions, which offers clearer physical interpretation in coordinate space. From these LFWFs, one can directly access partonic observables such as parton distribution functions (PDFs), relevant for high-energy processes [73, 74, 75]. Connections between rest-frame wave functions and LFWFs, particularly in the absence of a magnetic field, have been extensively utilized in previous studies (e.g., [76, 77, 78, 79, 80, 81, 82]), providing a solid foundation for the current work.
@@ II.1 Quark model in a magnetic field
where $r=\sqrt{z^{2}+r_{\perp}^{2}}$, $C$ is a constant shift, and the magnitudes of the confining, color Coulomb, and spin-spin potentials are characterized by the parameters, $\sigma$, $A$, and $\alpha$, respectively. We smear the spin-spin interaction as a Gaussian function with a parameter $\Lambda$ [85], for simplicity. The term $\Braket{\bm{S}_{q}\cdot\bm{S}_{\bar{q}}}$ yields the values of 1/41/41 / 4 and 3/4-3/4- 3 / 4 for the vector (spin-triplet) and pseudoscalar (spin-singlet) mesons, respectively.
@@
$\bra{10}-(\bm{\mu}_{1}+\bm{\mu}_{2})\cdot\bm{B}\ket{00}$ $=$ $-\frac{gqB}{4\mu},$ (8)
$\bra{00}-(\bm{\mu}_{1}+\bm{\mu}_{2})\cdot\bm{B}\ket{10}$ $=$ $-\frac{gqB}{4\mu},$ (9)
where the spin eigenstates are labeled by |$\ket{SS_{z}}$ with $S=0,1$ and $S_{z}=0,\pm 1$. Thus, only the longitudinal component |$\ket{10}$ of $S=1$ can mix with the $S=0$ state whereas the transverse components |$\ket{1\pm 1}$ do not mix with other components. Taking into account the off-diagonal interaction, we should solve a coupled-channel Schrödinger equation. We note that in the case of quarkonia, all the diagonal matrix elements, such as $\bra{00}-(\bm{\mu}_{1}+\bm{\mu}_{2})\cdot\bm{B}\ket{00}$, are vanishing: there is no hadronic Zeeman shift.
@@ II.2 Cylindrical Gaussian Expansion Method
where $c_{n}$ represents the expansion coefficient. The Gaussian basis function in position space is given by
@@
×$\times\mathrm{e}^{-ip_{z}z}\mathrm{e}^{-i\bm{r}_{\perp}\cdot\bm{p}_{\perp}}\phi^{\mathrm{CG}}_{n}(z,\bm{r}_{\perp}),$
@@
The normalization factors of the basis function, $N_{n}$ and $\tilde{N}_{n}$, are determined by $\innerproduct{\phi_{n}^{\mathrm{CG}}}{\phi_{n}^{\mathrm{CG}}}=1$.
@@
where the Hamiltonian matrix elements are given by $H_{q\bar{q},nm}=\bra{\phi_{n}^{\mathrm{CG}}}\hat{H}\ket{\phi_{m}^{\mathrm{CG}}}$, and the elements of the overlap matrix are $S_{nm}=\Braket{\phi_{n}^{\mathrm{CG}}|\phi_{m}^{\mathrm{CG}}}$. The Gaussian range parameters $\beta_{n}$ and $\gamma_{n}$ are determined by adopting a geometric progression [71, 72]
@@
$\innerproduct{\psi}{\psi}=\sum_{n,m}c_{n}^{*}S_{nm}c_{m}=1.$ (17)
Since the spin singlet |$\ket{00}$ and longitudinal component |$\ket{10}$ of the spin triplet mix with each other in the magnetic field, the wave function should take into account the coupled channel. The basis function is given as a linear combination of spin-0 and spin-1 components
@@ II.3 Light-front wave functions
After diagonalizing the Hamiltonian (6), the resulting wave functions are transformed into LFWFs [76, 77, 79, 80, 81, 82]. These LFWFs are expressed in terms of internal variables that are invariant under Lorentz boosts in the longitudinal direction, defined as
@@
The spin and orbital angular momentum part $\mathcal{R}^{JJ_{z}}_{\lambda_{q}\lambda_{\bar{q}}}$ of the meson LFWF is obtained through the interaction-independent Melosh transformation [90], which relates the instant-form spin states to the light-front helicity basis. Here $J$ and $J_{z}$ denote the total spin and its projection along the quantization axis, respectively, while $\lambda_{q}$ and $\lambda_{\bar{q}}$ are the light-front helicities of the quark and antiquark. This procedure assumes that the meson is described in terms of a non-interacting quark–antiquark basis, in line with the Bakamjian–Thomas construction [91, 92, 93, 94]. This procedure connects the spin-orbital structure of the ordinary rest-frame wave function, which carries definite quantum numbers $J^{PC}$, to its light-front representation in a consistent way [91, 77]. The effect of the external magnetic field is first included in the rest-frame wave function, by diagonalizing the Hamiltonian (6). We assume that these modifications are then carried over to the LFWFs via the Melosh transformation. Since our goal is to understand how the LFWFs are modified in the nonrelativistic quark model, this approach captures the dominant effects of the magnetic field such as spin-mixing and quark Landau levels.
For spin-orbital angular momentum part, it is convenient to work with the covariant form as [95, 96]
@@ II.4 Numerical Setup
$m_{Q}$ $\alpha$ [GeV] $\sqrt{\sigma}$ [GeV] $\Lambda$ [GeV2] $A$ $C$ [GeV]
$(c\bar{c})$ 1.784 0.4778 0.402 1.020 0.713 0.5693-0.5693- 0.5693
@@ III.1 Charmonia
The mass spectra of charmonia under strong magnetic fields are shown in Fig. 1. These results were previously discussed in Ref. [22]. Here, we recalculate the magnetic field dependence of the mass spectra using a denser set of data points. One crucial aspect is the optimization procedure for the range parameters, $[\beta_{1},\beta_{n_{\mathrm{max}}}]$ and $[\gamma_{1},\gamma_{n_{\mathrm{max}}}]$. Since the wave function becomes deformed into a cigar-like shape under strong magnetic fields, the optimal range parameters vary with $eB$. With improved optimization by Optim.jl [103] for each data point, we confirm the magnetic field dependence of the mass spectra reported in Ref. [22]. Although the nonrelativistic approximation works well when the magnetic field is smaller than the quark mass, we plot the mass spectra up to $eB=10$ GeV2 for theoretical exploration, where the difference from the relativistic approach can be useful to discuss the relativistic effect.
@@
[…]
@@ III.2 Bottomonia
The magnetic field dependence of the bottomonium mass spectra is shown in Fig. 7, where we recalculated with a denser set of data points compared with previous works [22]. The transverse vector states [$\Upsilon(1S,2S,3S)_{T}$] are displayed in the left panel, and the pseudoscalar [$\eta_{b}(1S,2S,3S)$] and longitudinal vector [$\Upsilon(1S,2S,3S)_{L}$] states, affected by mixing, are shown in the right panel. In the right panel, we compute the six lowest eigenstates below the $B\bar{B}$ threshold and label them as the 1st-6th states. This allows us to track the avoided crossings and evolution of the eigenstates in the presence of mixing induced by the $-\bm{\mu}_{i}\cdot\bm{B}$ interaction.
@@
Figure 13: LFWFs of the $J/\psi_{T}$ state at $eB=0$ and 1.01.01.0 GeV2 with a positive or negative correction (37). The longitudinal momentum fraction distribution with a positive (negative) correction becomes more squeezed (elongated).
Figure 14: PDFs of the $J/\psi_{T}$ states at various magnetic field strengths with a positive or negative correction (37) included. Up to $eB$=1 GeV2, the PDFs with the positive (negative) correction become slightly narrower (broader) as the magnetic field increases.
@@ III.3 Relativistic correction
[…]
@@
The impact of the additional $k_{z}^{2}$ term (37) on the mass spectrum is illustrated in Fig. 12. While this term slightly increases or decreases the mass of the $J/\psi_{T}$ depending on the sign, the linear trajectory remains unaffected. The corresponding LFWFs, shown in Fig. 13, with a positive (negative) correction are now more squeezed (elongated) in the longitudinal direction. One important implication is that the PDFs become slightly narrower or broader, as illustrated in Fig. 14. For comparison, in the light-front Hamiltonian approach [24], the PDFs for charmonium states corresponding to $\eta_{c}\to$ 1st and $J/\psi_{T}$ [^1] In Ref. [24], these states are labeled by a special quantum number $m_{PC}=-1$ and +1+1+ 1. at $eB=1\,\text{GeV}^{2}$ are slightly broader and narrower, respectively (see Fig. 9 of Ref. [24]). These behaviors may reflect complex relativistic dynamics but might be roughly interpreted as the $k_{z}^{2}$ correction (37). Furthermore, Fig. 15 shows the magnetic-field dependence of $\expectationvalue{\tfrac{1}{2}k_{\perp}^{2}}$ and $\expectationvalue{k_{z}^{2}}$. Notably, the inclusion of the $k_{z}^{2}$ correction with a plus sign reverses the behavior of $\expectationvalue{k_{z}^{2}}$, changing from an increasing to a decreasing trend with growing magnetic field strength. This reversal is directly reflected in the modification of the PDFs.
@@ Appendix A Relativistic corrections
$=m+\frac{p_{z}^{2}+2n|qB|}{2m}-\frac{(p_{z}^{2}+2n|qB|)^{2}}{8m^{3}}+\dots$ (40)
[…]
@@
The factor of $-p_{z}^{2}/2m^{2}$ may be regarded as a correction to the original energy shift of Landau levels.
@@ Appendix B Potential matrix elements
$\Braket{r_{\perp}^{2}k_{z}^{2}}$ $=-\matrixelement{\phi_{n}^{CG}}{r_{\perp}^{2}\frac{\partial^{2}}{\partial z^{2}}}{\phi_{m}^{CG}}$
@@
where $k_{z}^{2}$ corresponds to $-\partial^{2}/\partial z^{2}$ in coordinate space.

### v2

# Structure of heavy quarkonia in a strong magnetic field

@@ I Introduction
Recent theoretical advances have enabled comprehensive studies of these phenomena through multiple approaches. Lattice QCD simulations now provide first-principles calculations of hadron properties in magnetic fields [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19] ([20] for a review), while model-based frameworks offer complementary perspectives. This vibrant research area is evidenced by numerous recent reviews covering various aspects of quarkonia in extreme environments [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31].
@@
The study of heavy hadrons containing charm or bottom quarks under magnetic fields has been approached through several theoretical frameworks. Constituent quark models [32, 33, 34, 35, 36, 37, 38] and QCD sum rules [39, 40, 41, 42] have provided valuable insights, complemented by holographic QCD approaches [43, 44, 45, 46, 47] and the complex-potential picture in thermal medium [48, 49, 50, 51, 52]. Lattice QCD simulations have further advanced our understanding by investigating the heavy quark-antiquark potential in magnetic backgrounds [53, 54, 55]. Beyond heavy-quark systems, significant attention has been devoted to understanding magnetic field effects on hadrons composed of light (up, down, or strange) quarks. For light mesons, studies (e.g., [56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67]) have revealed characteristic modifications, while light baryons have been examined through several approaches [68, 69, 63, 70, 71, 72, 73, 74, 75, 76, 77].
@@
The constituent quark model remains a powerful framework for understanding static hadron properties, as detailed in a recent review [29]. Magnetic fields influence hadrons through three main mechanisms: (i) spin mixing via the $\bm{\mu}_{i}\cdot\bm{B}$ interaction, (ii) quark kinetic energy modifications induced by the $\bm{B}\times\bm{r}$ term, and (iii) anisotropic deformation of inter-quark potentials [78, 79, 80, 81], where lattice QCD simulations provide crucial constraints [53, 54, 55]. Previous investigations of quarkonia [35, 36] have systematically examined these effects, particularly spin mixing and energy shifts, employing potential models. These studies utilize the cylindrical Gaussian expansion method (CGEM) to solve the two-body Schrödinger equation, an adaptation of the conventional Gaussian expansion method (GEM) [82, 83] specifically designed for systems with cylindrical symmetry under magnetic fields.
@@
Building upon previous studies [36, 35], we aim to study light-front wave functions (LFWFs) of heavy quarkonia from nonrelativistic quark model solutions under strong magnetic fields and analyze their structural modifications. Unlike the direct light-front Hamiltonian diagonalization method proposed in Ref. [38], our approach establishes a connection between magnetically-modified LFWFs and the rest-frame wave functions, which offers clearer physical interpretation in coordinate space. From these LFWFs, one can directly access partonic observables such as parton distribution functions (PDFs), relevant for high-energy processes [84, 85, 86]. Connections between rest-frame wave functions and LFWFs, particularly in the absence of a magnetic field, have been extensively utilized in previous studies (e.g., [87, 88, 89, 90, 91, 92, 93, 94]), providing a solid foundation for the current work.
@@ II.1 Quark model in a magnetic field
where $r=\sqrt{z^{2}+r_{\perp}^{2}}$, $C$ is a constant shift, and the magnitudes of the confining, color Coulomb, and spin-spin potentials are characterized by the parameters, $\sigma$, $A$, and $\alpha$, respectively. We smear the spin-spin interaction as a Gaussian function with a parameter $\Lambda$ [97], for simplicity. The term $\Braket{\bm{S}_{q}\cdot\bm{S}_{\bar{q}}}$ yields the values of $1/4$ and 3/4-3/4 for the vector (spin-triplet) and pseudoscalar (spin-singlet) mesons, respectively.
@@
[…]
@@ II.2 Cylindrical Gaussian Expansion Method
where $c_{n}$ represents the expansion coefficient. The Gaussian basis function in position space is given by [35, 36]
@@
$\times\mathrm{e}^{-ip_{z}z}\mathrm{e}^{-i\bm{r}_{\perp}\cdot\bm{p}_{\perp}}\phi^{\mathrm{CG}}_{n}(z,\bm{r}_{\perp}),$
@@
The normalization factors of the basis function, $N_{n}$ and $\tilde{N}_{n}$, are determined by $\braket{\phi_{n}^{\mathrm{CG}}|\phi_{n}^{\mathrm{CG}}}=1$.
@@
[…]
@@
$\braket{\psi|\psi}=\sum_{n,m}c_{n}^{*}S_{nm}c_{m}=1.$ (17)
[…]
@@ II.3 Light-front wave functions
After diagonalizing the Hamiltonian (6), the resulting wave functions are transformed into LFWFs [87, 88, 89, 90, 91, 92, 93, 94]. These LFWFs are expressed in terms of internal variables that are invariant under Lorentz boosts in the longitudinal direction, defined as
@@
The spin and orbital angular momentum part $\mathcal{R}^{JJ_{z}}_{\lambda_{q}\lambda_{\bar{q}}}$ of the meson LFWF is obtained [102, 103, 104] through the interaction-independent Melosh transformation [105], which relates the instant-form spin states to the light-front helicity basis. Here $J$ and $J_{z}$ denote the total spin and its projection along the quantization axis, respectively, while $\lambda_{q}$ and $\lambda_{\bar{q}}$ are the light-front helicities of the quark and antiquark. This procedure assumes that the meson is described in terms of a non-interacting quark–antiquark basis, in line with the Bakamjian–Thomas construction [106, 107, 108, 109]. This procedure connects the spin-orbital structure of the ordinary rest-frame wave function, which carries definite quantum numbers $J^{PC}$, to its light-front representation in a consistent way. The effect of the external magnetic field is first included in the rest-frame wave function, by diagonalizing the Hamiltonian (6). We assume that these modifications are then carried over to the LFWFs via the Melosh transformation. Since our goal is to understand how the LFWFs are modified in the nonrelativistic quark model, this approach captures the dominant effects of the magnetic field such as spin-mixing and quark Landau levels.
For the spin-orbital angular momentum part, it is convenient to work with the covariant form as [110, 111]
@@ II.4 Numerical Setup
$m_{Q}$ $\alpha$ [GeV] $\sqrt{\sigma}$ [GeV] $\Lambda$ [GeV2] AA $C$ [GeV]
$(c\bar{c})$ 1.784 0.4778 0.402 1.020 0.713 0.5693-0.5693
@@ III.1 Charmonia
The mass spectra of charmonia under strong magnetic fields are shown in Fig. 1. These results were previously discussed in Ref. [36]. Here, we recalculate the magnetic field dependence of the mass spectra using a denser set of data points. One crucial aspect is the optimization procedure for the range parameters, $\beta_{1},\beta_{n_{\mathrm{max}}}]$ and $\gamma_{1},\gamma_{n_{\mathrm{max}}}]$. Since the wave function becomes deformed into a cigar-like shape under strong magnetic fields, the optimal range parameters vary with $eB$. With improved optimization by Optim.jl [118] for each data point, we confirm the magnetic field dependence of the mass spectra reported in Ref. [36]. Although the nonrelativistic approximation works well when the magnetic field is smaller than the quark mass, we plot the mass spectra up to $eB=10$ GeV2 for theoretical exploration, where the difference from the relativistic approach can be useful to discuss the relativistic effect.
@@
In contrast, the mass shifts of the pseudoscalar states ($\eta_{c}(1S)$ and $\eta_{c}(2S)$) and longitudinal vector states ($J/\psi_{L}$ and $\psi(2S)_{L}$) are shown in the right panel of Fig. 1. In this case, the pseudoscalar-vector mixing occurs due to the $\bm{\mu}_{i}\cdot\bm{B}$ term, so that we label the resulting states as 1st, 2nd, 3rd, and 4th. The 1st and 2nd states decrease and increase, respectively, as a result of level repulsion with increasing magnetic field. Around eB=1eB=133 GeV2 we find avoided crossings such as 2nd-3rd and 3rd-4th. We also note that in the strong magnetic field region, the level structure becomes simpler.
@@ III.2 Bottomonia
The magnetic field dependence of the bottomonium mass spectra is shown in Fig. 7, where we recalculated with a denser set of data points compared with previous works [36]. The transverse vector states $\Upsilon(1S,2S,3S)_{T}$] are displayed in the left panel, and the pseudoscalar $\eta_{b}(1S,2S,3S)$] and longitudinal vector $\Upsilon(1S,2S,3S)_{L}$] states, affected by mixing, are shown in the right panel. In the right panel, we compute the six lowest eigenstates below the $B\bar{B}$ threshold and label them as the 1st-6th states. This allows us to track the avoided crossings and evolution of the eigenstates in the presence of mixing induced by the $\bm{\mu}_{i}\cdot\bm{B}$ interaction.
@@
Figure 13: LFWFs of the $J/\psi_{T}$ state at $eB=0$ and $1.0$ GeV2 with a positive or negative correction (37). The longitudinal momentum fraction distribution with a positive (negative) correction becomes more squeezed (elongated).
Figure 14: PDFs of the $J/\psi_{T}$ states at various magnetic field strengths with a positive or negative correction (37) included. Up to eBeB=1 GeV2, the PDFs with the positive (negative) correction become slightly narrower (broader) as the magnetic field increases.
@@ III.3 Relativistic correction
In this work, we phenomenologically introduce a $k_{z}$-dependent Landau level term as a relativistic correction (i.e., a higher-order term in the $1/m$ expansion):
@@
The impact of the additional $k_{z}^{2}$ term (37) on the mass spectrum is illustrated in Fig. 12. While this term slightly increases or decreases the mass of the $J/\psi_{T}$ depending on the sign, the linear trajectory remains unaffected. The corresponding LFWFs, shown in Fig. 13, with a positive (negative) correction are now more squeezed (elongated) in the longitudinal direction. One important implication is that the PDFs become slightly narrower or broader, as illustrated in Fig. 14. For comparison, in the light-front Hamiltonian approach [38], the PDFs for charmonium states corresponding to $\eta_{c}\to$ 1st and $J/\psi_{T}$ [^1] In Ref. [38], these states are labeled by a special quantum number $m_{PC}=-1$ and +1+1. at $eB=1\,\text{GeV}^{2}$ are slightly broader and narrower, respectively (see Fig. 9 of Ref. [38]). These behaviors may reflect complex relativistic dynamics but might be roughly interpreted as the $k_{z}^{2}$ correction (37). Furthermore, Fig. 15 shows the magnetic-field dependence of $\expectationvalue{\tfrac{1}{2}k_{\perp}^{2}}$ and $\expectationvalue{k_{z}^{2}}$. Notably, the inclusion of the $k_{z}^{2}$ correction with a plus sign reverses the behavior of $\expectationvalue{k_{z}^{2}}$, changing from an increasing to a decreasing trend with growing magnetic field strength. This reversal is directly reflected in the modification of the PDFs.
@@ Appendix A Relativistic corrections
$=m+\frac{p_{z}^{2}+2n|qB|}{2m}-\frac{(p_{z}^{2}+2n|qB|)^{2}}{8m^{3}}+\dots.$ (40)
In the last form, the first and second terms are consistent with the mass, the kinetic energy, and the nonrelativistic Landau levels with the cyclotron frequency $|qB|/m$, which are used in the usual nonrelativistic framework. On the other hand, the third term is the higher-order corrections proportional to $1/m^{3}$, which includes the cross term between $p_{z}^{2}$ and $2n|qB|$. Thus, the momentum $p_{z}$ and the Landau level effect can be correlated as a higher-order relativistic correction. The energy shift from this term is
@@
The factor of $p_{z}^{2}/2m^{2}$ may be regarded as a correction to the original energy shift of Landau levels.
@@ Appendix B Potential matrix elements
[…]
@@
where $k_{z}^{2}$ corresponds to $\partial^{2}/\partial z^{2}$ in coordinate space.